OpenStudy (anonymous):

Help! plz!! Evaluate if a=2/3, b=1/6 Give your answer in form of an fraction: a+b/a-b

OpenStudy (anonymous):

i know it's 1 but how do you get it

OpenStudy (anonymous):

\[\frac{\frac{2}{3}+\frac{1}{6}}{\frac{2}{3}-\frac{1}{6}}\] multiply top and bottom by 6 to clear the fractions

OpenStudy (anonymous):

\[\frac{\frac{2}{3}+\frac{1}{6}}{\frac{2}{3}-\frac{1}{6}}\times \frac{6}{6}\] \[=\frac{4+1}{4-1}\] which is not 1

OpenStudy (anonymous):

as a matter of fact it cannot be one, because the only way a fraction is one is if the top and bottom are the same, and \(a+b\neq a-b\)

OpenStudy (anonymous):

that's what i thought too

OpenStudy (anonymous):

thx